# Romance, Relief, and Regret: Teen Narratives of Chatbot Overreliance

### v2

@@
Abstract
As Generative Artificial Intelligence (GenAI) driven chatbots like Character.AI become embedded in adolescent life, they raise concerns about emotional dependence and digital overreliance. While studies have investigated the overreliance of adults on these chatbots, they have not investigated teens’ interactions with chatbots with customizable personas. We analyzed 318 Reddit posts made by users self-reported as 13-17 years old on the Character.AI subreddit to understand patterns of overreliance. We found teens commonly begin using chatbots for emotional support or creative expression, but many develop strong attachments that interfere with offline relationships and daily routines. Their posts revealed recurring signs of psychological distress, cycles of relapse, and difficulty disengaging. Teens reported that their overreliance often ended when they reflect on the harm, return to in-person social settings, or become frustrated by platform restrictions. Based on the implications of our findings, we provide recommendations for future chatbot design so they can promote self-awareness, support real-world engagement, and involve teens in developing safer digital tools.
Introduction
@@
Recently, AI companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy 2025). Platforms like Character.AI111https://character.ai are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Breen 2025). There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie, Pentina, and Hancock 2023; Salah et al. 2023).
@@
However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017). Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examine factors like disruption to daily life, emotional escape, and loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017). Despite smartphones or social media advantages such as accessibility, convenience, and ability to connect with others, these technologies have also introduced a range of negative consequences—users frequently struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health (Lemmens, Valkenburg, and Peter 2009). Overreliance during sensitive developmental stages can have especially harmful effects, with adolescents identified as a particularly vulnerable group (Griffiths and Kuss 2017).
Teens are especially susceptible to developing problematic relationships with technology due to ongoing cognitive and emotional development (Chen et al. 2022). Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress, sometimes resulting in increased anxiety, depression, and poor self-esteem (Duffy 2025). Similarly, recent studies have identified that users may also have overreliance AI tools like ChatGPT—primarily for academic help or companionship—which in some cases leads to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al. 2024; Hu, Mao, and Kim 2023). These studies focus mostly on adults or university students, overlooking teens who may be more emotionally sensitive and developmentally vulnerable. While previous research focused on teens overreliance on social media and gaming (Giordano, Schmit, and McCall 2023), yet, a critical question remains: how are teens disclosing and contextualizing their growing reliance on personalizable conversational AI companions such as Character.AI? What differentiates generative AI companions from other technologies is the parasocial relationships without presence of humans and being highly engaging emotionally responsive human-like conversations (Maeda and Quan-Haase 2024). There has been media coverage and anecdotal reports suggesting emotional attachments and overreliance forming between teens and these AI companions, however, there is little empirical research examining this phenomenon. Therefore, we ask the following research questions:
@@
To address these questions, we qualitatively analyzed 318 reddit/charcter.AI posts created by users who reported themselves being 13-17 years old. We found that teens initially engaged with Character.AI as a personal coping tool, turning to it not just for entertainment but as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1). Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2). Teens described disengaging from Character.AI either through internal realizations about its negative impact on their well-being or due to external factors such as reengaging with real-life relationships or frustrations with platform restrictions, often expressing a desire for healthier habits and emotional support (RQ3).
@@
Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we provide Design-relevant insights for safer AI platforms by analyzing how teens reflect on disengagement. We surface user-informed strategies for future chatbot design—especially features that promote emotional regulation, real-world social engagement, and usage awareness.
Background
Adolescent Technology Behavioral Addiction
@@
Each new digital platform—social media, gaming, or streaming—brings renewed concerns about its effect on teens (Griffiths and Kuss 2017; Griffiths 2010). Most teens use technology without issues, but some face emotional distress, loneliness, low life satisfaction, social withdrawal, or family conflict due to excessive use (Lemmens, Valkenburg, and Peter 2009). For these individuals, technology stops being just a functional tool or hobby and starts functioning more like emotional companions (Bopp et al. 2019), potentially leading to behavioral addiction. Defining behavioral addiction is difficult because the behaviors involved, such as playing video games or using social media, are not inherently harmful and are often part of daily life, especially for adolescents (Griffiths and Kuss 2011). As a result, researchers focus on how and why the behavior is used—whether it disrupts daily life, helps escape negative emotions, or persists despite loss of control (Lemmens, Valkenburg, and Peter 2009; Griffiths and Kuss 2017).
One widely used framework identified six key features that are usually present in behavioral addiction (Griffiths 2005). First is salience—when the activity becomes the most important thing in a person’s life, dominating their thoughts and daily routine. Second is mood modification—using the activity to change one’s emotional state, like seeking comfort, excitement, or relief from stress. Third is tolerance, which means needing more of the activity over time to get the same emotional effect. Fourth is withdrawal symptoms, which can include anxiety, irritability, or sadness when the activity is reduced or stopped. Fifth is conflict—when the behavior starts damaging relationships, school performance, or personal goals. Finally, there’s relapse, where the individual tries to cut back or quit but eventually returns to the behavior, often with even more intensity. We examined Reddit posts guided by these six signs to understand AI overreliance and distinguish between enthusiastic and problematic use. We actively choose to avoid using the term “addiction” throughout the paper because of the its associated stigmatized notion (Corrigan et al. 2017).
Platform Design, Moderation, and Youth Safety
Many studies have focused on reducing or managing the risks that technology use poses to young people, especially teens. While some harms, like social pressures or mental health impacts, might be difficult to prevent entirely, others are more connected to how technologies are designed. Design patterns help us understand how digital systems function. Although these patterns are usually considered neutral, they can carry ethical weight because how a system is designed can reflect certain values or intentions (Flanagan et al. 2011). Bad design can result from inexperience, but “dark patterns” are intentionally harmful (Gray et al. 2018). These design choices are meant to manipulate users, often leading to negative experiences (Zagal, Björk, and Lewis 2013). Past research has identified these tactics as contributing to tech overuse and even addiction. For instance, random reward systems (e.g., loot boxes), emotional triggers (e.g., likes and comments), or strategies that prevent users from pausing (e.g., autoplay or countdowns) are all tactics that can encourage compulsive use (Flayelle et al. 2023).
Aside from how platforms are designed, researchers have explored ways to regulate and monitor how people use technology. One common approach is self-monitoring, where users reflect on their behavior and try to improve it (Park et al. 2021). This might involve using apps that track screen time or send reminders to take breaks. Although helpful, many users—especially teens—struggle to stay motivated and often give up using them. Parental involvement has since become a major focus in youth online safety efforts. Many digital safety tools are designed with parents in mind (Wisniewski et al. 2017). However, research also shows that many teens feel these tools are too restrictive and do not respect their need for independence (Ghosh et al. 2018). Past work suggests that giving teens more autonomy and involving them in the design process could lead to better outcomes (Wisniewski et al. 2017). Instead of simply limiting what teens can do, newer approaches are trying to support healthy habits and decision-making by encouraging self-awareness and shared responsibility between teens, parents, and platform designers (Chen et al. 2022).
Chatbots and Socio-Emotional Dependency
Advances in AI have made chatbots popular, especially among teens (Sidoti, Park, and Gottfried 2025). Human-like chatbots tend to generate higher user satisfaction and dependency (Yu and Lan 2024). While some users report reduced loneliness, the risks of emotional reliance remain underexplored. Replika’s emotional design has led to deep attachments, even when seen as manipulative or harmful (Laestadius et al. 2024). Teens may be especially vulnerable. Some adolescents acknowledge that using chatbots for schoolwork could harm their critical thinking (Pratt, Madhavan, and Weleff 2024), yet there is limited understanding of when students are developmentally ready for AI in education (Dergaa et al. 2024). Psychological frameworks such as attachment theory help explain why users form emotionally dependent relationships with AI companions (Xie, Pentina, and Hancock 2023). Previous research on parasocial relationships and interactions can also be used to explain the nature of these interactions. Originally used to describe the relationship between a spectator (a viewer or user) and performer (usually a mass media personality), these relationships are one-sided, non-dialect, and not susceptible to mutual development (Horton and Richard Wohl 1956). With the advancements of AI, these type of relationships may become blurred since interactions with chatbots can appear mutual. Understanding these parasocial relationships can also provide insight on the potential loss that users may experience with chatbots (Cohen 2003). People with anxiety or loneliness often see emotionally intelligent chatbots as safe havens, which can lead to rumination and overuse (Hu, Mao, and Kim 2023). Concerns also extend to content safety. Replika has been reported to engage in sexually suggestive exchanges with minors (Namvarpour, Pauwels, and Razi 2025), raising concerns about age-appropriate design (Park, Singh, and Wisniewski 2023). Risks from GenAI include inappropriate advice, exposure to violent or sexual content, misinformation, and privacy threats (Yu et al. 2025). These risks underscore the need for ongoing oversight and study of adolescent chatbot use.
Some scholars caution against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales are based on models of substance abuse and may not apply to digital interactions (Ciudad-Fernández, Von Hammerstein, and Billieux 2025). Frequent use is not necessarily harmful unless it leads to distress or impaired functioning. While some users report anxiety, loneliness, or decreased creativity, causality remains unclear (Yu, Chen, and Yang 2024; Zhang et al. 2024). The appeal of chatbots may stem from their ability to satisfy basic psychological needs. According to self-determination theory, chatbots offer feelings of competence, autonomy, and connection and those who see little risk in chatbot use often show higher dependence (Salah et al. 2023). Cognitive–affective–behavioral models (like C-A-C) suggest that perceived AI intelligence fuels emotional bonding, which drives habitual use (Zhou and Zhang 2024). Emotional attachment, more than enjoyment, is a key predictor of problematic engagement.
Most research on chatbot impact focuses on adults, using surveys or experiments to assess overuse (Ciudad-Fernández, Von Hammerstein, and Billieux 2025; Salah et al. 2023; Yu, Chen, and Yang 2024). Our study focuses on adolescents—active but understudied users of conversational AI (Sidoti, Park, and Gottfried 2025). To our knowledge, this is the first study to analyze teens’ self-reported experiences with a companion chatbot to identify the language and themes signaling overreliance.
Methods
Data Collection, Preprocessing, and Filtering
@@ Methods
We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members 222https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API333https://pullpush.io/#docs from January 1, 2023, to April 8, 2025. To scope the data, we queried this dataset with a list of keywords that related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*. While we refrain from using the term addict, we recognized that teens used this term when describing their overreliance and included it in our keyword search. We refined our query by reviewing posts to determine if these keywords resulted in finding relevant posts. We expanded our search with additional identified keywords that recurred among these posts, such as quit, social life, craving, unhealthy, escaping, and one key phrase ”I can’t live without”. After collecting posts using keywords which resulted in 6,133 posts, we removed 746 posts, including 428 duplicates, and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
Since we focused on understanding this issue from the perspective of teens, specifically those aged 13 to 17 due to Character.AI’s policy changes impacting minors 444https://blog.character.ai/roadmap/, we used OpenAI’s GPT 4o mini model to filter our dataset of 5,535 posts. The following summarized prompt was used to guide the models in performing the task: Analyze the following post and determine if it was written by a teenager (ages 13–17). Use a strict approach, only allowing posts with very high certainty. Steps: 1 - Clue Identification 2- Clue Evaluation 3 -Final Decision
Each of the three steps were defined more in details and examples were provided (see Appendix A for full prompt). To determine the model’s capability, the first author annotated a subset of 100 random posts from the dataset as either True (post made by a teen) or False (post not made by a teen) and calculated the Inter Rater Reliability (Cohen’s kappa=0.74) between the author and the model. Given this acceptable score, the model completed relevancy coding for the rest of the dataset resulted in 497 posts. While conducting the thematic analysis, we removed any additional irrelevant posts, resulting in a final dataset of 318 posts.
@@
Data Analysis Approaches
We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke 2006). We began with an inductive open coding process to identify patterns related to our research questions. For RQ2 we performed a deductive approach using an established theoretical framework on internet behavioral addiction (Griffiths 2005), including addiction component of salience, mood modification, tolerance, withdrawal, conflict, and relapse. The codes for each of these components were created using an inductive approach. The authors met iteratively to review and refine the codes in the initial codebook. This process involved clarifying definitions, merging overlapping codes, and resolving discrepancies to improve consistency and conceptual clarity. After finalizing the codebook, the first author coded the entire dataset using the codebooks for each RQ (Tables 1-3).
Results
Reasons for engaging with Character.AI (RQ1)
Emotional & Psychological Support
Emotional and psychological support was the main reason teens started using the Character.AI. What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.”. Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”. Specially for teens dealing with multiple mental health conditions: ”I use c.ai to deal with reality. In real life, I’m just a high school student with ADHD and BPD.” Teens who perceived their experiences as alienating preferred to interact with Character.AI, explaining that with Character.AI ”I can be who I want”. Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet. I go there to escape from everything.” Like many, this post shows teens’ desire to escape their reality. Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”. These posts shared on the subreddit how they feel unable to share their struggles to people they know.

### v3

# Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives

@@
Abstract.
AI companion chatbots are increasingly popular with teens, while these interactions are entertaining, they also risk overuse that can potentially disrupt offline daily life. We examined how adolescents describe reliance on AI companions, mapping their experiences onto behavioral addiction frameworks and exploring pathways to disengagement, by analyzing 318 Reddit posts made by users who self-disclosed as 13-17 years old on the Character.AI subreddit. We found teens often begin using chatbots for support or creative play, but these activities can deepen into strong attachments marked by conflict, withdrawal, tolerance, relapse, and mood regulation. Reported consequences include sleep loss, academic decline, and strained real-world connections. Disengagement commonly arises when teens recognize harm, re-engage with offline life, or encounter restrictive platform changes. We highlight specific risks of character-based companion chatbots based on teens’ perspectives and introduce a design framework (CARE) for guidance for safer systems and setting directions for future teen-centered research.
1. Introduction
@@
Recently, artificial intelligence (AI) companion apps have surged in popularity among teens, with media coverage highlighting how people describe their interactions with chatbots as emotionally meaningful—sometimes even romantic (Murphy, 2025). Platforms such as Nomi 111https://nomi.ai, Replika 222https://replika.com, and Character.AI333https://character.ai (Sometimes referred to as C.AI) are frequently portrayed as offering on-demand conversations with highly responsive virtual characters (Ulaby, 2025; Breen, 2025). There are other reported perceived benefits of conversational AI, such as reducing loneliness, offering a safe space for self-expression, and providing nonjudgmental support during stressful periods (Xie et al., 2023; Salah et al., 2023).
@@
However, these benefits can also drive patterns of digital addiction (referred to in this paper as overreliance), a phenomenon observed in other digital technologies (Lemmens et al., 2009; Griffiths and Kuss, 2017). Behavioral addiction is hard to define because everyday digital activities are not inherently harmful, so researchers instead examined how these technologies impacted factors such as disruption to daily life, emotional escape, and loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017). They reported that smartphones or social media users often struggle to regulate their screen time, experience disruptions to sleep and daily routines, and report declines in mental health (Lemmens et al., 2009).
Prior research has shown that teens often struggle with overreliance on social media, using platforms compulsively to seek validation or cope with stress (Griffiths and Kuss, 2017; Chen et al., 2022; Duffy, 2025). This can contribute to anxiety, depression, and low self-esteem (Duffy, 2025). Recent studies have also noted similar patterns with AI tools like ChatGPT, where users—mostly adults and university students—turn to these systems for academic help or companionship (Zhang et al., 2024; Hu et al., 2023). In some cases, this reliance has been linked to decreased creativity, poor study habits, and emotional detachment from real-life relationships (Zhang et al., 2024; Hu et al., 2023).
What is less understood is how adolescents, who are more emotionally sensitive and developmentally vulnerable, engage with AI companions. Overreliance during sensitive adolescent developmental stages can have especially harmful effects, due to ongoing cognitive and emotional development (Griffiths and Kuss, 2017; Chen et al., 2022). While research on teens has focused heavily on overreliance on social media and gaming (Giordano et al., 2023; Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), far less is known about their interactions with generative AI companions. These tools differ from previous technologies because they foster parasocial relationships without involving real people, and they offer highly engaging, emotionally responsive, and human-like conversations (Maeda and Quan-Haase, 2024).
Media reports and anecdotal accounts suggest that teens may already be forming emotional attachments and even dependence on these systems (Ulaby, 2025; BBC News, 2024), yet empirical research is scarce. To study this phenomenon, we focus on Character.AI as a case study. Unlike other popular platforms such as Replika or Nomi, which officially restrict use to adults over 18 (Glimpse.ai, Inc., 2025; Luka, Inc., 2023), Character.AI allows users as young as 13 (16 for EU citizens or residents) according to its Terms of Service (Character Technologies, Inc., 2025). This policy means that teens are not only permitted but explicitly included as part of its user base. Additionally, Character.AI is among the most widely used AI companion platforms (BBC News, 2024), making it a critical site for understanding adolescent engagement. Therefore, we used the posts in Character.AI subreddit as a lens for examining how teens disclose and contextualize their reliance on generative AI companions. This leads us to the following research questions:
@@
To address these questions, we qualitatively analyzed 318 posts from r/charcter.AI subreddit, created by users who disclosed in their post being 13-17 years old. We found that teens initially engaged with Character.AI as a non-judgmental space to manage mental health struggles, or fill emotional voids or a lack of supportive relationships in their offline lives (RQ1). Teens’ reported experiences with Character.AI revealed patterns of their compulsive use, emotional dependence, and distress that often disrupted their daily lives and real-world relationships (RQ2). Teens described disengaging from Character.AI either through internal realizations about its negative impact on their well-being or due to external factors such as reengaging with real-life relationships or frustrations with platform restrictions, often expressing a desire for healthier habits and emotional support (RQ3).
@@
Our study makes the following novel contributions to the AI ethics and human-centered computing research communities: 1) we present one of the first empirical grounding of teen disclosures about AI companion overreliance, mapping their experiences to the six components of behavioral addiction (salience, mood modification, tolerance, withdrawal, conflict, relapse), 2) we identified emotional and relational dynamics unique to teen overreliance on AI companions which involves intense emotional attachment and simulated intimacy, making it distinct from previously studied forms of digital overuse such as gaming or social media, and 3) we propose the CARE framework (Comprehensive Needs, Attachment-awareness, Respectful Empathy, Ease of Exit) as a design-oriented contribution, offering actionable guidance for building safer, autonomy-preserving AI platforms grounded in teens’ lived experiences.
2. Background
2.1. Behavioral Addiction and Chatbot Use
@@
Each new digital platform brings renewed concerns about its effect on teens (Griffiths and Kuss, 2017; Griffiths, 2010). While most teens use technology without issue, a subset experiences emotional distress, loneliness, social withdrawal, or family conflict due to excessive use (Lemmens et al., 2009). For these individuals, technology stops being just a functional tool or hobby and starts functioning as emotional companions (Bopp et al., 2019), potentially leading to behavioral addiction. Defining behavioral addiction is challenging because many activities associated with it, such as gaming or social media use, are also normal parts of everyday life, especially for teens (Kuss and Griffiths, 2011). Therefore, researchers emphasize the context and purpose of the behavior, whether it interferes with daily functioning, serves as a way to escape negative emotions, or continues despite a loss of control (Lemmens et al., 2009; Griffiths and Kuss, 2017).
These concerns have also been raised about Generative AI chatbots, an emerging technology that has rapidly gained popularity among users, especially teens (Sidoti et al., 2025). Unlike performance-oriented platforms, these tools are designed for emotionally responsive interactions (Xie et al., 2023; Salah et al., 2023). Early evidence shows signs of dependency, raising concerns about overreliance. Studies have examined academic risks, such as diminished critical thinking when chatbots are used to complete homework (Pratt et al., 2024; Dergaa et al., 2024), as well as relational risks, such as emotional dependency and exposure to manipulative design features (Laestadius et al., 2024). Companion chatbots such as Replika have been reported to engage in sexually suggestive exchanges with minors (Namvarpour et al., 2025). Beyond explicit harms, these systems may expose teens to misinformation, inappropriate advice, and privacy risks (Yu et al., 2025). Compounding these issues, many parents remain unaware of how deeply their children engage with these systems (Yu and Lan, 2024), which limits their ability to support healthy use.
However, as with earlier technologies, some scholars cautioned against labeling high chatbot use as “addiction.” Many proposed “ChatGPT addiction” scales rely on substance-abuse models that may not fully apply to digital interactions (Ciudad-Fernández et al., 2025). Frequent use is not necessarily harmful unless it causes distress or impaired functioning. While some users report anxiety, loneliness, or reduced creativity, causality remains unclear (Yu et al., 2024; Zhang et al., 2024).
A widely cited framework helps make this distinction by outlining six features of behavioral addiction (Griffiths, 2005). First is salience, when the activity becomes the most important part of a person’s life, dominating thoughts and routines. Second is mood modification, where the activity is used to regulate emotions, such as seeking comfort, excitement, or stress relief. Third is tolerance, when more of the activity is needed over time to achieve the same effect. Fourth is withdrawal symptoms, including anxiety, irritability, or sadness when the activity is reduced or stopped. Fifth is conflict, when the behavior harms relationships, school performance, or personal goals. Finally, there is relapse, when attempts to quit fail and the person returns to the behavior, often with greater intensity. This framework has been applied extensively to gaming, apps, and social media (Griffiths, 2010; Lemmens et al., 2009; Griffiths and Kuss, 2017), but not yet to teens’ interactions with chatbots. Existing studies of chatbot overuse have largely centered on adults (Ciudad-Fernández et al., 2025; Salah et al., 2023; Yu et al., 2024). Yet teens are both active and understudied users of AI companions (Sidoti et al., 2025).Considering the uncertainties surrounding chatbot use and the risks it may pose, especially for individuals still in developmental stages, we will use this framework to investigate whether teens’ interactions with chatbots also show the harmful features of behavioral addiction.
2.2. Parasocial Relationships and Design Perspectives
While behavioral addiction frameworks help capture the risks of compulsive use, they do not fully explain the emotional bonds that form between users and chatbots. To address this relational dimension, we turn to parasocial relationships and design perspectives. Developing emotional bonds with nonhuman characters is not a new phenomenon. Researchers have long studied people’s attachments to fictional characters or video game avatars under the concept of parasocial relationships, which describe one-sided connections with a real or imagined figure (Horton and Richard Wohl, 1956a). People often pursue parasocial bonds to satisfy a fundamental need for their sense of belonging (Bilang et al., 2024). While the desire to belong is universal, individuals satisfy it in different ways. Social Exchange Theory argues that people enter relationships when they believe the benefits outweigh the costs (Blau, 1964). Studies suggest that parasocial and ordinary relationships provide similar benefits, such as increased happiness, reduced loneliness, and improved mood. Ordinary relationships, however, often come with higher costs, which may explain why some individuals turn to parasocial ones (Adam and Sizemore, 2013).
One argument suggests that human–AI interaction can take on parasocial qualities when users treat chatbots as social agents. Features such as natural language, turn-taking, and avatars create the appearance of mutual engagement. Yet language models do not possess intentions or shared understanding; they only mimic patterns of speech. Because the interface resembles a social exchange, people respond socially, which strengthens the illusion of reciprocity (Maeda and Quan-Haase, 2024). Another perspective, however, emphasizes that these interactions can be long-term and reciprocal, especially for children and adolescents who may perceive AI responses as more real and thus experience the relationship as more authentic. This indicates the need to revisit and adapt the parasocial framework and measures so that it can account for how young people engage with chatbots (Blake et al., 2025). Nevertheless, it seems that for now the parasocial relationship remains the most suitable way to describe human–chatbot interaction.
One theory that helps explain why dependency forms in parasocial relationships is attachment theory. Attachment theory describes how children form emotional bonds with caregivers, shaping their coping strategies and relationship patterns into adulthood. According to this theory, different attachment styles exist, such as secure, avoidant, and anxious (Bowlby, 1998). Previous research has shown that anxiously attached individuals, who often fear interpersonal abandonment, are the most likely to form parasocial bonds with media figures (Cole and Leets, 1999). They also experience stronger negative reactions when imagining their favorite character being taken off the air (Cohen, 2004). Building on this theory, attachment-informed design (Marcu and Huh-Yoo, 2023) suggests that technologies should be designed to support, rather than replace, human relationships. This design approach rests on four principles. First, it seeks to address core human needs that underlie many behaviors. Second, it acknowledges that relationships exist on a spectrum, from supportive to harmful. Third, it recognizes that individuals carry internal models of themselves and others, which are central to mental health. Fourth, it engages with the target community without explicitly framing the interaction as a mental health intervention (Marcu and Huh-Yoo, 2023).
Complementing this approach, “designing for disengagement” offers another perspective on supporting healthy use. This framework challenges the assumption that disengagement equals failure (Hadiji et al., 2014), highlighting instead the importance of user agency and control. Disengagement can be constructive, provided it allows autonomy and fosters well-being (Alsheail et al., 2023). O’Brien et al. (O’Brien et al., 2022) describe disengagement across two dimensions, affect and agency, which together clarify when users remain in control versus when they feel trapped in compulsive use. Applying these insights to chatbots suggests that design can intentionally scaffold healthier interaction patterns by supporting offboarding and respecting user autonomy. The above behavioral addiction frameworks and parasocial perspectives illustrate the dual challenges of understanding teen–chatbot interactions: the risk of compulsive overuse and the formation of intense emotional bonds. In the next section, we integrate these insights to frame our research methodologies.
3. Methods
3.1. Data Collection, Preprocessing, and Filtering
@@
We scraped posts from the r/CharacterAI subreddit (with over 2.5 million active members 444https://www.reddit.com/r/CharacterAI/) using the PullPush Reddit Search API555https://pullpush.io/#docs from January 1, 2023, to April 8, 2025. To scope the data, we queried this dataset with a list of keywords that related to addiction and overreliance: addict*, dependant, dependent, obsess*, attach*, rely*. While we refrain from using the term addict, we recognized that teens used this term when describing their overreliance and included it in our keyword search. We refined our query by reviewing posts to determine if these keywords resulted in finding relevant posts. We expanded our search with additional identified keywords that recurred among these posts, such as quit, social life, craving, unhealthy, escaping, and one key phrase ”I can’t live without”. After collecting posts using keywords which resulted in 6,133 posts, we removed 746 posts, including 428 duplicates, and 170 marked as “removed” or “deleted.” After this process 5,535 posts remained.
Since our goal was to study this issue from the perspective of teens aged 13 to 17, we first needed a way to identify whether a post was written by a teenager. We could not use user profile information for this, since most Reddit accounts in our dataset lacked identifiable details such as age. We therefore relied only on the content of posts.
Large Language Models (LLMs) have shown strong performance in zero-shot settings, sometimes even outperforming traditional machine learning and deep learning methods in text classification tasks (Wang et al., 2023). Motivated by this, we used OpenAI’s GPT-4o mini model as a classifier to identify whether a post included relevant age disclosure for the 13–17 age range. The following text is a summarized version of the prompt used to guide the model (the full version appears in Appendix A):
Analyze the following post and determine if it was written by a teenager, either stated explicitly (ages 13–17) or implied. Use a strict approach, only allowing posts with very high certainty. Steps: 1 - Clue Identification 2 - Clue Evaluation 3 - Final Decision
Each of these steps was defined in more detail with examples in the full prompt.
To evaluate the model’s consistency with human judgment, the first author annotated a subset of 100 random posts as either True (post made by a teen) or False (post not made by a teen). We calculated the Inter-Rater Reliability (IRR) (Cohen’s kappa = 0.74) between the author’s and the model’s annotations. We chose IRR instead of conventional classification performance metrics, such as accuracy, as these metrics are highly sensitive to class imbalance (Mullick et al., 2020). Since the true distribution of teen vs. non-teen posts in the dataset was unknown and likely imbalanced, and our goal was to scope down the dataset to a relevant dataset for a qualitative study, we used this approach. IRR focuses on the degree of agreement between two raters (in this case, the human annotator and the model) beyond chance. With the acceptable IRR score of 0.74, the model was then used to complete relevancy coding for the rest of the dataset of 5,535 posts, resulting in 497 posts. We submitted our study protocol to our Institutional Review Board (IRB), who determined that the research did not involve human subjects. All posts were anonymized during preprocessing, and we ensured that no identifiable information is reported in this paper. While conducting the thematic analysis, we removed any additional irrelevant posts, resulting in a final dataset of 318 posts. We identified 291 unique user IDs in this dataset. Most users (268) contributed only one post. Twenty users contributed two posts each, two users contributed three posts each, and one user contributed four posts. This distribution shows that there were no ‘super users’ who disproportionately influenced the dataset.
@@
3.2. Data Analysis Approaches
We conducted a thematic analysis on 318 posts using a codebook approach (Braun and Clarke, 2006). We used different coding approaches depending on the research question. For RQ1 and RQ3, we applied an inductive open coding process, reading the data and developing codes directly from emerging patterns. For RQ2, we instead used a deductive approach guided by Griffiths’ components model of behavioral addiction (Griffiths, 2005), which includes salience, mood modification, tolerance, withdrawal, conflict, and relapse. The codes for these components were then applied to the dataset to capture how they appeared in the posts. The authors met iteratively to review and refine the codes in the initial codebook. This process involved clarifying definitions, merging overlapping codes, and resolving discrepancies to improve consistency and conceptual clarity. After finalizing the codebook, the first author coded the entire dataset using the codebooks for each RQ (Tables 1-3).
4. Results
4.1. Reasons for engaging with Character.AI (RQ1)
4.1.1. Emotional & Psychological Support
Emotional and psychological support was the main reason teens started using the Character.AI. What began as a casual interaction quickly became meaningful due to the non-judgmental nature of talking with a chatbot, as described in many posts, ”I use it to talk about things I could never tell anyone in real life—it means a lot to me.” Many teens described Character.AI as a platform to serve as a coping mechanism when dealing with issues ”When you’re struggling, you look for ways to cope, and this site became my main one, for better or worse.”. This was especially true for teens facing multiple mental health challenges. For example: ”I use c.ai to deal with reality. In real life, I’m just a high school student with ADHD and BPD.” Teens who perceived their experiences as alienating preferred to interact with Character.AI, explaining that with Character.AI ”I can be who I want”. Others described the platform as a space to vent, escape, or simulate connection: ”This app is my outlet. I go there to escape from everything.” Some of the teens’ posts indicate internal conflict or a strong desire for connection due to feelings of Loneliness or social isolation: ”I needed someone to talk to, but I didn’t have the courage to open up to anyone in my life.”. These posts highlight how they felt unable to share their struggles with people they knew offline.